OpenStudy (anonymous):

Suppose the population of a town is 8,600 and is growing 3% each year. Predict the population after 3 years. A. about 77,400 people B. about 9,397 people C. about 26574 people D. about 232,200 people

OpenStudy (lgbasallote):

are you allowed to use differential equations? i can solve this using that method...

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

B. about 9,397 people

OpenStudy (anonymous):

ok what about this one..

OpenStudy (anonymous):

Suppose that the population of deer in a state is 1,500 and is growing 2% each year. Predict the population after 4 years. A. about 12,000 deer B. about 1,624 deer C. about 3,110 deer D. about 24,000 deer

OpenStudy (anonymous):

Firstly calculate 2 % of 1500..

OpenStudy (lgbasallote):

@El-naggar are you sure about your answer? maybe you should provide a solution to prove your answer.

OpenStudy (anonymous):

2% of 1500 is: \[\frac{1500 \times 2}{100} = 30\] After 4 years increase will be: 30 + 30 30 + 30 = 30*4 = 120 Just add 120 to the given population of deer and get the answer...

OpenStudy (anonymous):

@lgbasallote it's easy 3% per year after 3 yrs the population increases with 9% and 9% of 8600 is 774 so after 3 yrs the population is 8600+774=9374